OpenStudy (anonymous):

Solve 5x2 - 7x + 2 = 0 by completing the square. Will give medals. Please work step by step

OpenStudy (anonymous):

ok for completing the square we must always bring out the common factor before completing the square. \[5(x ^{2}\frac{-7}{5}x + \frac{2}{5}) = 0\] We should always look at coefficient of the middle term. so the coefficient is -7/5 of the middle term. now we want to divide that by 2 and square it. \[(\frac{-7}{5})^{2}=\frac{49}{25}\] now we have to add and minus it onto the equation. \[5{(x ^{2}-\frac{7}{5}+\frac{49}{25}-\frac{49}{25}+\frac{2}{5})}=0\] see if u know what to do from here.